### v3

# Performance Analysis of Indoor VLC Network with Secure Downlink NOMA for Body Blockage Model

Abstract
In this work, we investigate the performance of indoor visible light communication (VLC) networks based on power domain non-orthogonal multiple access (NOMA) for mobile devices, where multiple legitimate users are equipped with photodiodes (PDs). We propose a body blockage model for both the legitimate users and eavesdropper to address scenarios where the communication links from transmitting light-emitting diodes (LEDs) to receiving devices are blocked by the bodies of all parties. Furthermore, in order to improve the secrecy without requiring knowledge of the channel state information (CSI) of the eavesdropper, we propose a novel LED arrangement that reduces the overlapping areas covered by different LED units supporting distinct users. We also suggest two LED transmission strategies, i.e., simple and smart LED linking, and compare their performance with the conventional broadcasting in terms of transmission sum rate and secrecy sum rate. Through computer simulations, we demonstrate the superiority of our proposed strategies to the conventional approach.
Index Terms:
Body blockage model, non-orthogonal multiple access (NOMA), passive eavesdropper, physical layer security (PLS), visible light communication (VLC).
I Introduction
Visible light communication (VLC) is a promising technology for the beyond fifth-generation (B5G) and sixth-generation (6G) networks [1]. Operating in the visible light spectrum, VLC utilizes the light emitted for both illumination and data transmission. With its high potential data rate, VLC has garnered significant interest [2]. Implemented through intensity modulation (IM) of light-emitting diodes (LEDs), VLC leverages existing infrastructure for illumination. It serves as an intriguing complement to radio frequency (RF) communication, particularly due to the need for strict RF radiation monitoring.
Non-orthogonal multiple access (NOMA) has gained attention for its high spectral efficiency compared to the conventional orthogonal multiple access (OMA) techniques [3, 4, 5]. Unlike OMA, NOMA allocates the same time-frequency resources to multiple users, maximizing the limited frequency spectrum utilization [6]. This work focuses on power domain NOMA, employing multi-user superposition coding at the transmitter and successive interference cancellation (SIC) at the receivers [7]. Power allocation algorithms for NOMA have been investigated in previous studies, e.g., [8, 9].
In addition, physical layer security (PLS) facilitates secure communication by leveraging wireless channel randomness, including fading, noise, and interference [10]. PLS aims to utilize the stochastic properties of wireless channels to provide secure communication. It ensures secrecy when the channel of the legitimate user surpasses that of the eavesdropper, making eavesdropping-resistant information transmission feasible without the need for symmetric keys or complex encryption/decryption algorithms [11, 12, 13, 14]. Also, PLS with VLC-enabled NOMA has been investigated in [15, 16, 17, 18]. Specifically, in [15], a precise expression of secure outage probability was derived for a NOMA-enabled multi-user multiple-input single-output (MISO) VLC system. In [16], secure transmission of MISO with NOMA under VLC, considering both known and unknown channel state information (CSI) of eavesdroppers, was investigated. In [17], secrecy performance of a NOMA-enabled VLC system in the presence of an active eavesdropper was demonstrated. In [18], secrecy performance is enhanced in the NOMA-enabled VLC system by the pseudo user, which interrupts the SIC process of the eavesdropper.
I-A Literature Review and Motivation
In recent years, a variety of signal processing approaches has been investigated for VLC-aided PLS systems such as beamforming [19, 20, 21], precoding [22, 23], artificial optical jamming [24], and polar code-based secure coding [25]. Regarding mathematical tools, stochastic geometry theory is used to derive the average secrecy capacity of the system while the users are randomly distributed across the room [26]. Convex optimization theory can be used to enhance the secrecy performance of VLC systems that are subject to various illumination constraints, e.g., the LED optical power constraint [27], the peak power constraint to reduce eye damage [19], and the light energy harvesting and dimming control constraint [20]. However, most of the studies on PLS for VLC networks have focused on defending against external malicious eavesdroppers [23, 26, 27], some of which addressing the transmission of confidential information where users are unaware of information not intended for them [22]. Based on the properties of VLC channels, these studies considered applications of beamforming and jamming for enhancing the secrecy in the physical layer with high computational complexity. In practice, however, the position of the passive eavesdropper is unknown since they do not respond to transmission mechanisms. Furthermore, it is still highly complex for broadcasting systems to implement beamforming schemes in VLC transmission. With respect to LED arrangements, most works adopt a square lattice alignment to improve illuminance [28]. However, in broadcasting VLC, such an arrangement may degrade communication reliability since interference may occur in the overlapping areas served by multiple LED devices with each transmitting their own signal.
Furthermore, previous researches on indoor VLC systems ignore the effects of body blockage from users, with few exceptions such as [29, 30, 31]. In [29], the authors investigated the cylinder model of user bodies without considering the angles. Reference [30] investigates the use of reconfigurable intelligent surfaces (RIS) in order to avoid body blockage in VLC system. In [31], the modeling of body blockage has been investigated by dividing the orientation angle into several parts, each of which was assessed for blockage.
I-B Contributions
Inspired by the above-mentioned state-of-the-art studies, our work focuses on VLC-enabled NOMA-based PLS systems without any knowledge of the eavesdropper. The main contributions111 The preliminary version of this work was presented at the 2023 IEEE Wireless Communications and Networking Conference (WCNC) in [32]. In the present manuscript, we additionally propose a new strategy (i.e., smart LED linking strategy) as well as a new power allocation scheme based on an estimated maximum sum rate in order to improve the secrecy performance. of this work are summarized as follows:
  • In the aforementioned works (e.g., [31, 33]), the modeling of body blockage has only been partially addressed. However, the modeling methods of body blockage in these works are not universally applicable, i.e., these models are based on the statistical methods to approximate the channel state from LED sources to PD receivers. Therefore, we propose a geometrical-based general blockage model in the presence of a passive eavesdropper. This model accounts for scenarios where the light path from the LED to the PD may be blocked by user bodies, with blockage events determined through geometric calculations.
  • A new power allocation scheme is proposed based on NOMA to maximize the transmission sum rate of legitimate users, considering only their locations and statistical radiation angles. The system, however, is unaware of the devices’ azimuth and polar angles, which are important for detecting whether the light is blocked. Furthermore, two power allocation schemes are compared through computer simulations.
  • In order to decrease the overlapping area of different LED units, we propose a new LED arrangement, where LEDs are placed in the form of an equilateral triangle lattice. Furthermore, we introduce two transmission strategies for this LED arrangement, called simple and smart LED linking strategies, so as to investigate the trade-off between the transmission and secrecy sum rates.
I-C Paper Organization
The rest of this paper is organized as follows. Section II describes the system model including the channel model as well as the body blockage model considered throughout this work, and Section III describes the NOMA signal models and performance metrics. In Section IV, the proposed transmission strategies of secure NOMA in VLC are described. Simulation results are presented and discussed in Section V. Finally, Section VI concludes this work.
Notations
TABLE I: Notations
Notation Description
$[x]^{+}$ An operator is defined as $[x]^{+}=\max(x,0)$.
${\mathbb{E}}[\cdot]$ Expectation operator.
${\cal N}(\mu,\sigma^{2})$ The circular-symmetric Gaussian random distribution with mean $\mu$ and variance $\sigma^{2}$.
[…]
$\mathbbm{1}\left[{\cal X}\right]$ A binary indicator function for an event ${\cal X}$ that is equal to $1$ if the event ${\cal X}$ occurs and 00 otherwise.
$\|{\bm{x}}\|$ Euclidean norm of the vector ${\bm{x}}=(x_{1},x_{2},\ldots,x_{M})$, i.e., $\|{\bm{x}}\|=\sqrt{\sum_{m=1}^{M}x_{m}^{2}}$
The notations used in this paper are listed in Table I.
II System and Channel Model
In this section, we describe our VLC system model as well as VLC channel model.
II-A VLC System Model
Figure 1: An illustration of the system model. Cyan dots represent LED transmitters, $D_{1},\cdots,D_{K}$ denote the positions of PDs of the legitimate users, and $D_{E}$ is the position of the eavesdropper.
As shown in Fig. 1, we consider an indoor downlink scenario in the room size of length $L$, width $W$, and height $Z$, where $N$ LED sources serve $K$ legitimate users with the index set ${\cal K}=\{1,2,\ldots,K\}$, as well as an eavesdropper denoted by $D_{E}$. We consider that each of the legitimate users and the eavesdropper has a single PD receiver. We assume that the system is not aware of the location of the eavesdropper. Moreover, we assume that all the bodies of users as well as the eavesdropper are modeled as cylinders with radius $r$ as in [35, 36, 31]. Also, let $S_{n}=(x_{S_{n}},y_{S_{n}},Z)$ denote the position corresponding to the $n$th LED source in the Cartesian coordinates for $n\in{\cal I}\triangleq\{1,2,\ldots,N\}$. We define $D_{k}=(x_{D_{k}},y_{D_{k}},z_{D_{k}})$ with $k\in{\cal K}$ as the positions representing the corresponding PD of legitimate users, and $D_{E}=(x_{E},y_{E},z_{E})$ representing the corresponding PD of the eavesdropper. Furthermore, we define ${\cal K}^{\prime}\triangleq{\cal K}\cup E$ as the set of indices corresponding to all the PD receivers in the room.
In addition, we assume that the height of the device $z_{D_{k}}$ and eavesdropper $z_{E}$ are defined as the constant ratio $\nu\in(0,1)$ of the height of the users, i.e., $z_{D_{k}}=\nu{\sf h}_{D_{k}}$ and $z_{D_{E}}=\nu{\sf h}_{E}$.
II-B Channel Model
In this subsection, we define the VLC channel between an LED-PD transceiver pair with emphasis on the effect of the users’ bodies.
II-B1 VLC Channel
Figure 2: An illustration of the VLC path between the $n$th LED and the $k$th PD, where $\theta_{k,n}$ is the radiation angle between $S_{n}$ and $D_{k}$, $\theta_{1/2}$ is the half illuminance angle of the LED transmitter, $\psi_{k,n}$ is the incidence angle of link between $S_{n}$ and $D_{k}$, $\Psi$ is the received field of view (FoV) of the PD receiver, and ${\bm{n}}_{k}$ is the normal direction of the photosensitive surface of the $k$th PD.
Figure 3: The geometry of visible light communications with $S_{n}$, $D_{k}$ and $U_{k}$. $S_{n}^{\prime}$ is the projection point of $S_{n}$ onto the plane that contains the top base of cylinder, ${\bm{n}}_{k}$ is the normal direction of the photosensitive surface of $D_{k}$ with its projection at XOY plane is given by ${\bm{n}}_{k}$, $\omega_{k}$ is the azimuth angle between ${\bm{n}}_{k}$ and $x$-axis, $\lambda_{k}$ is the polar angle of $D_{k}$, $\theta_{k,n}$ is the radiation angle from $S_{n}$ to $D_{k}$, $\psi_{k,n}$ is the incidence angle of the light beam from $S_{n}$ to $D_{k}$, and $\phi_{k,n}$ is the azimuth angle between the vector $\overrightarrow{S_{n}U_{k}}$ and $y$-axis.
In Fig. 2, we illustrate the geometric relationship between LED $S_{n}$ and PD $D_{k}$ with the normal direction of its photosensitive surface ${\bm{n}}_{k}$. Most of the existing studies on VLC systems consider the line-of-sight (LoS) component in VLC channel models and ignore the non-line-of-sight (NLoS) component due to a huge diffuse component gap between LoS and NLoS components [37, 38, 39]. Subsequently, assuming that the considered LEDs have Lambertian emission patterns, the channel gain between $S_{n}$ and $D_{k}$ with $k\in{\cal K}^{\prime}$ is expressed as [37]
$h_{k,n}\left(\theta_{k,n},\psi_{k,n}\middle|\Psi\right)$
$=A\frac{(m+1)R_{\mathrm{PD}}}{2\pi}\cos^{m}(\theta_{k,n})\frac{\cos(\psi_{k,n})}{d_{k,n}^{2}}g\left(\psi_{k,n}\middle|\Psi\right)$
$\quad\cdot\prod_{l\in{\cal K}^{\prime}}\mathbbm{1}\left[{\cal E}(n,k,l)\right],$ (1)
where the notations introduced in Section II-B1 are listed in Table II.
TABLE II: Parameter Definitions in Equation Section II-B1
Notation Description
$A$ The effective area of the photosensitive surface of PD receivers.
$\theta_{k,n}$ The radiation angle from $S_{n}$ to $D_{k}$.
$\psi_{k,n}$ The incidence angle from $S_{n}$ to $D_{k}$.
$\Psi$ The received field of view (FoV) of PD receivers.
$m$ The order of the Lambertian emission with half illuminance angle of LED transmitter $\theta_{1/2}$, defined as $m=\frac{-\ln(2)}{\ln\left(\cos\left(\theta_{1/2}\right)\right)}$.
$d_{n,k}$ The Euclidean distance between $S_{n}$ and $D_{k}$.
$g\left(\psi_{n,k}\middle|\Psi\right)$ The gain of the optical concentrator with reflective index $\eta\in(1,2)$, expressed as $g\left(\psi_{k,n}\middle|\Psi\right)=\begin{cases}\frac{\eta^{2}}{\sin^{2}(\Psi)},&\leavevmode\nobreak\ \mbox{if}\leavevmode\nobreak\ \leavevmode\nobreak\ 0\leq\psi_{k,n}\leq\Psi,\\ 0,&\leavevmode\nobreak\ \mbox{if}\leavevmode\nobreak\ \leavevmode\nobreak\ \psi_{k,n}>\Psi.\end{cases}$
In Section II-B1, ${\cal E}(n,k,l)$ represents the event that the link from the $n$th source $S_{n}$ to the $k$th user $D_{k}$ is not blocked by the body of the $l$th user $D_{l}$. In order to facilitate the understanding of these geometric parameters, we depict the geometric relations of $S_{n}$, $D_{k}$, and $U_{k}$ in Fig. 3, where $U_{k}=(x_{U_{k}},y_{U_{k}},{\sf h}_{D_{k}})$ is the position of the top center of the cylinder for $k\in{\cal K}^{\prime}$. The vector from $D_{k}$ to $U_{k}$, $\overrightarrow{D_{k}U_{k}}\triangleq\left(x_{U_{k}}-x_{D_{k}},y_{U_{k}}-y_{D_{k}},{\sf h}_{D_{k}}-z_{D_{k}}\right)$, can be expressed as
$\overrightarrow{D_{k}U_{k}}$ $=\overrightarrow{S_{n}U_{k}}-\overrightarrow{S_{n}D_{k}}$
$=\left(\sqrt{l_{d}^{2}-z_{D_{k}}^{2}}\cos\omega_{k},\sqrt{l_{d}^{2}-z_{D_{k}}^{2}}\sin\omega_{k},\Delta z_{D_{k}}\right),$ (2)
where $l_{d}$ is the distance between the center of cylinder and the PD of its mobile device, $\Delta z_{D_{k}}\triangleq{\sf h}_{D_{k}}-z_{D_{k}}$, and $\omega_{k}\in[-\pi,\pi)$ is the azimuth angle between the line from $S_{n}$ to $D_{k}$ and $y$-axis, i.e., the angle between the XOY projection of the direction vector ${\bm{n}}_{k}$ and $x$-axis [40]. Moreover, the incidence angle $\psi_{k,n}$ in Section II-B1 can be obtained by the positions of $S_{n}$ and $D_{k}$, azimuth angle $\omega_{k}$, and polar angle $\lambda_{k}\in[0,\frac{\pi}{2}]$, i.e., [33]
$\cos\psi_{k,n}$
$=\left(\frac{x_{S_{n}}-x_{D_{k}}}{\left\|\overrightarrow{D_{k}S_{n}}\right\|}\right)\sin\lambda_{k}\cos\omega_{k}$
$\quad+\left(\frac{y_{S_{n}}-y_{D_{k}}}{\left\|\overrightarrow{D_{k}S_{n}}\right\|}\right)\sin\lambda_{k}\sin\omega_{k}+\left(\frac{Z-z_{D_{k}}}{\left\|\overrightarrow{D_{k}S_{n}}\right\|}\right)\cos\lambda_{k},$ (3)
[…]
II-B2 Body Blockage Model
(a) Three-dimensional (3D) schematic diagram.
(b) Two-dimensional (2D) plane (XOY) schematic diagram.
Figure 4: A mathematical explanation of the body blockage model in this work. The $n$th LED light is positioned at $S_{n}(x_{S_{n}},y_{S_{n}},Z)$, while the PD receiver for the $k$th user is located at $D_{k}(x_{D_{k}},y_{D_{k}},z_{D_{k}})$. A cylinder, who is modeled as the receiver holder of $D_{l}$ with $l\in{\cal K}^{\prime}$, is centered at $U_{l}(x_{U_{l}},y_{U_{l}},{\sf h}_{D_{l}})$ on its top plane, with an azimuth angle $\phi_{l,n}$ between itself and the $n$th LED source. The light vector ${\bm{a}}_{k,n}$ projects onto the XOY plane as $\tilde{\bm{a}}_{k,n}$. The plane ${\cal P}_{l,n}$, which passes through the point $U_{l}$, is perpendicular to the vector $\tilde{\bm{a}}_{k,n}$. Additionally, the vector ${\bm{a}}_{k,n}$ intersects with the planes ${\cal Q}_{l}$ and ${\cal P}_{l,n}$ at the points $Q_{l,n}$ and $P_{l,n}$, respectively.
The presence of the body of users as well as the eavesdropper inside an indoor environment may certainly affect the channel characteristics and the communication performance of a VLC system. In fact, the channel gain of the PD devices is affected not only by its geometric parameters but also by the shadows of the human bodies. Subsequently, we focus on the event ${\cal E}(n,{k},{l})$ in Section II-B1. The indicator function $\mathbbm{1}\left[{\cal E}(n,k,l)\right]$ plays a pivotal role in determining whether this event occurs, returning 1 if it does and 0 otherwise. In what follows, we describe the role of this function in detail. As illustrated in Fig. 4, in order to simplify notations, let us define ${\bm{a}}_{k,n}=\overrightarrow{S_{n}D_{k}}=(x_{D_{k}}-x_{S_{n}},y_{D_{k}}-y_{S_{n}},z_{D_{k}}-Z)$ as a vector of light from $S_{n}$ to $D_{k}$ with the projection on the XOY plane $\tilde{\bm{a}}_{k,n}=(x_{D_{k}}-x_{S_{n}},y_{D_{k}}-y_{S_{n}},0)$. Moreover, we define two planes ${\cal Q}_{l}$ and ${\cal P}_{l,n}$, where ${\cal Q}_{l}$ contains the point $U_{l}$ with its unit normal vector given by ${\bm{q}}=(0,0,1)$, and ${\cal P}_{l,n}$ contains the point $U_{l}$ with its unit normal vector given by ${\bm{p}}_{l,n}=\alpha\tilde{\bm{a}}_{k,n}$ for a real-valued factor $\alpha$. The projection area of the vertical cylinder with the four vertices on the plane ${\cal P}_{l,n}$ is a rectangle, with their coordinates given by
$P_{l,n}^{(1)}$ $=\left(x_{U_{l}}+r\cos\phi_{l,n},y_{U_{l}}-r\sin\phi_{l,n},{\sf h}_{D_{l}}\right),$ (4a)
$P_{l,n}^{(2)}$ $=\left(x_{U_{l}}-r\cos\phi_{l,n},y_{U_{l}}+r\sin\phi_{l,n},{\sf h}_{D_{l}}\right),$ (4b)
$P_{l,n}^{(3)}$ $=\left(x_{U_{l}}-r\cos\phi_{l,n},y_{U_{l}}+r\sin\phi_{l,n},0\right),$ (4c)
$P_{l,n}^{(4)}$ $=\left(x_{U_{l}}+r\cos\phi_{l,n},y_{U_{l}}-r\sin\phi_{l,n},0\right),$ (4d)
where $\phi_{l,n}$ is the azimuth angle between $S_{n}$ and $U_{l}$, defined as
$\phi_{l,n}=\arctan\frac{\left|y_{S_{n}}-y_{U_{l}}\right|}{\left|x_{S_{n}}-x_{U_{l}}\right|}.$ (5)
Next, the intersection point of a line and a plane satisfies the following Lemma:
Lemma 1.
Figure 5: Illustration for the relation of intersection point of vectors and plane ${\cal W}$ in Lemma 1. Points $N$ and $P$ are on plane ${\cal W}$, ${\bm{v}}_{s}$ is the normal vector to plane ${\cal W}$, and ${\bm{v}}_{t}$ is the vector have same direction as $\overrightarrow{MP}$ (i.e., ${\bm{v}}_{t}=a\overrightarrow{MP}$ with real factor $a$).
Suppose that the vector $\mathbf{v}_{t}=(x_{t},y_{t},z_{t})$ in Euclidean space passes through the point $M(x_{m},y_{m},z_{m})$ and the point $P(x_{p},y_{p},z_{p})$ on the plane ${\cal W}$ as illustrated in Fig. 5. If the plane ${\cal W}$ also includes the point $N(x_{n},y_{n},z_{n})$ with one of its normal vectors $\mathbf{v}_{s}=(x_{s},y_{s},z_{s})$, the vector $\overrightarrow{MP}$ should satisfy
$\overrightarrow{MP}=\frac{\overrightarrow{MN}\cdot\mathbf{v}_{s}}{\mathbf{v}_{s}\cdot\mathbf{v}_{t}}\mathbf{v}_{t},$ (6)
where the inner product of ${\bm{v}}_{s}$ and ${\bm{v}}_{t}$ satisfies
${\mathbf{v}_{s}\cdot\mathbf{v}_{t}}\neq 0.$ (7)
Proof.
See Appendix A. ∎
We define $Q_{l,n}^{({\rm inter})}$ as the intersection of vector ${\bm{a}}_{k,n}$ and plane ${\cal Q}_{l}$, and $P_{l,n}^{({\rm inter})}$ as the intersection of vector ${\bm{a}}_{k,n}$ and plane ${\cal P}_{l,n}$. Their coordinates can be derived from Lemma 1 as
$\overrightarrow{S_{n}Q_{l,n}^{({\rm inter})}}$ $=\frac{\overrightarrow{S_{n}U_{l}}\cdot{\bm{a}}_{k,n}}{{\bm{a}}_{k,n}\cdot{\bm{q}}}{\bm{q}},$ (8)
$\overrightarrow{S_{n}P_{l,n}^{({\rm inter})}}$ $=\frac{\overrightarrow{S_{n}U_{l}}\cdot{\bm{a}}_{k,n}}{{\bm{a}}_{k,n}\cdot{\bm{p}}_{l,n}}{\bm{p}}_{l,n}.$ (9)
Therefore, the event ${\cal E}_{k}(n,{k},{l})$ in Section II-B1 can be determined by the following proposition:
Proposition 1.
The path ${\bm{a}}_{k,n}$ is blocked by the cylinder body of user $U_{l}$, (i.e., the event ${\cal E}(n,{k},{l})$ in Section II-B1 is false), if either of the following two conditions is satisfied:
  1. 1.
    The intersection point $Q_{l,n}^{({\rm inter})}$ should lie in the top base of cylinder ${\cal Q}_{l}$, i.e., it should satisfy that
    $\|Q_{l,n}^{({\rm inter})}-U_{l}\|\leq r.$ (10)
  2. 2.
    The intersection point $P_{l,n}^{({\rm inter})}$ should be in the area of the rectangle formed by the points $P_{l,n}^{(1)}$, $P_{l,n}^{(2)}$, $P_{l,n}^{(3)}$, and $P_{l,n}^{(4)}$, i.e., it should satisfy both of the following inequalities
$\left(\overrightarrow{P_{l,n}^{(1)}P_{l,n}^{(2)}}\times\overrightarrow{P_{l,n}^{(1)}P_{l,n}^{({\rm inter})}}\right)\cdot\left(\overrightarrow{P_{l,n}^{(3)}P_{l,n}^{(4)}}\times\overrightarrow{P_{l,n}^{(3)}P_{l,n}^{({\rm inter})}}\right)$ $\geq 0,$
$\left(\overrightarrow{P_{l,n}^{(2)}P_{l,n}^{(3)}}\times\overrightarrow{P_{l,n}^{(2)}P_{l,n}^{({\rm inter})}}\right)\cdot\left(\overrightarrow{P_{l,n}^{(4)}P_{l,n}^{(1)}}\times\overrightarrow{P_{l,n}^{(4)}P_{l,n}^{({\rm inter})}}\right)$ $\geq 0.$ (11)
Otherwise, it is not blocked by the cylinder (i.e., the event ${\cal E}(n,{k},{l})$ in Section II-B1 is true).
III NOMA Signal Models and Performance Metrics
In this section, we describe the NOMA signal model for our proposed VLC transmission system and introduce performance metrics of the transmission and secrecy sum rate.
III-A NOMA Signal Model
We assume that the transmission signals by LED transmitters are the direct current (DC) biased signals [28]. For convenience, we define a new set $\mathcal{K}_{n}\subset\mathcal{K}$, which represents the indices of users that receive the signal transmitted from the $n$th LED $S_{n}$. The transmitted symbol from the $n$th LED, denoted by $x_{n}\in{\mathbb{R}}$, is a Gaussian signal with a fixed bias current $I_{\mathrm{DC}}\in{\mathbb{R}}_{+}$ that is applied for the purpose of illumination. Following the definition in [28, eq. (2)], we have333 For the signal power of VLC transmission, $I_{\rm DC}$ is a large constant value. This assumption is commonly used in the studies related to VLC, e.g., [28, 37, 41].
$x_{n}=\sum_{k\in{\cal K}_{n}}\sqrt{P_{s}\beta_{k}({\cal K}_{n})}s_{n,k}+I_{\mathrm{DC}},$ (12)
where $P_{s}$ represents the total transmission power at each LED and $s_{n,k}$ is the modulated message signal intended for $D_{k}$ by $S_{n}$, which is assumed to be a Gaussian random variable with zero-mean and unit variance444 Provided that $I_{\rm DC}$ is sufficiently large, the probability of negative $x_{n}$ could be ignored. , i.e., $s_{n,k}\sim{\cal N}(0,1)$. Furthermore, $\beta_{k}({\cal K}_{n})$ represents the allocated power ratio for $D_{k}$ and fulfills that
$\sum_{k\in{\cal K}_{n}}\beta_{k}({\cal K}_{n})=1.$ (13)
Therefore, after removing the DC bias, the received signal by $D_{k}$ for $k\in{\cal K}^{\prime}$ can be formulated as
$y_{k}$ $=\sum_{n=1}^{N}{h}_{k,n}\left(\theta_{k,n},\psi_{k,n}\middle|\Psi\right)\sum_{k_{n}\in{\cal K}_{n}}\sqrt{P_{s}\beta_{k_{n}}({\cal K}_{n})}x_{n}+n_{k},$ (14)
where $h_{k,n}\left(\theta_{k,n},\psi_{k,n}\middle|\Psi\right)$ is the channel coefficient from $S_{n}$ to $D_{k}$ defined in Section II-B1 and $n_{k}$ is the real additive white Gaussian noise (AWGN) with variance $\sigma_{N}^{2}$ at $D_{k}$, i.e., $n_{k}\sim{\cal N}(0,\sigma_{N}^{2})$.
III-B Transmission and Secrecy Sum Rate
We assume that each device can decode its information with perfect SIC [42], i.e., there is no residual interference after the SIC process. The device first detects the symbol with the highest received power and then starts the SIC process by treating the other symbols as additional Gaussian noise terms. The transmission rate555 The units of transmission rate as well as secrecy rate are defined as bit/Hz/sec [43]. of the $k$th PD $D_{k}$ for $k\in{\cal K}_{n}$, defined as $R_{D_{k}}$, is evaluated as [43]
$R_{D_{k}}\left(\theta_{k,n},\psi_{k,n}\right)$ $=\frac{1}{2}\log_{2}\left(1+\gamma_{k}\left(\theta_{k,n},\psi_{k,n},{\cal K}_{n}\right)\right),$ (15)
where the scaling factor $\frac{1}{2}$ is due to the baseband signal transmission and $\gamma_{k}(\theta_{k,n},\psi_{k,n},{\cal K}_{n})$ is the received signal-to-interference-plus-noise ratio (SINR) at the $k$th user (or eavesdropper) under the assumption of the perfect SIC, defined in Eq. 17 located at the top of the next page, with $|{\cal K}_{n}|$ representing the cardinality of the set ${\cal K}_{n}$. We define a new set for all the indices of LED light signal received by the $k$th user (or eavesdropper) in the set ${\cal K}_{n}$ as $\mathcal{I}_{k}=\{n\leavevmode\nobreak\ |\leavevmode\nobreak\ k\in\mathcal{K}_{n},n\in\mathcal{I}\}$, e.g., if the $k$th user received the signals transmitted by LED sources with indices 1 and 2, then ${\cal I}_{k}=\{1,2\}$. Also, $\mathcal{I}_{k}^{\tilde{n}}$ is defined in the same manner as $\mathcal{I}_{k}$ with an additional condition $\tilde{n}\neq n$, i.e., $\mathcal{I}_{k}^{\tilde{n}}=\{n\leavevmode\nobreak\ |\leavevmode\nobreak\ k\in\mathcal{K}_{n},n\in\mathcal{I},n\neq\tilde{n}\}$.
The eavesdropper wiretaps information sent from $S_{n}$ to $D_{k}$ during the downlink transmission period and employs the same decoding process as that performed at legitimate users. The transmission rate, which can be wiretapped by the eavesdropper from the data intended for the $k$th user via the eavesdropper’s channel, defined as $R_{E,D_{k}}$, is evaluated as
$R_{E,D_{k}}\left(\theta_{E,n},\psi_{E,n}\right)=\frac{1}{2}\log_{2}\left(1+\gamma_{k}(\theta_{E,n},\psi_{E,n},{\cal K}_{n})\right),$ (16)
where $\gamma_{k}(\theta_{E,n},\psi_{E,n},{\cal K}_{n})$ denotes the SINR of the $k$th user signal wiretapped by the eavesdropper, expressed by Eq. 17 located at the top of this page.
$\gamma_{k}(\theta_{k,n},\psi_{k,n},{\cal K}_{n})$ $=\displaystyle\frac{\overbrace{\max_{{\cal K}_{n}}\left[\left(\sum_{n\in{\cal I}_{k}}h_{k,n}\left(\theta_{k,n},\psi_{k,n}\middle|\Psi\right)\right)^{2}\beta_{k}({\cal K}_{n})\right]}^{{\rm Intended\leavevmode\nobreak\ signal\leavevmode\nobreak\ for\leavevmode\nobreak\ the\leavevmode\nobreak\ }{\it k}{\rm th\leavevmode\nobreak\ user\leavevmode\nobreak\ from\leavevmode\nobreak\ }{\it{\cal I}_{k}}}}{\underbrace{\max_{{\cal K}_{n}}\left[\left(\sum_{n\in{\cal I}_{k}}h_{k,n}\left(\theta_{k,n},\psi_{k,n}\middle|\Psi\right)\right)^{2}\sum_{\overline{k}={k+1}}^{\left|{\cal K}_{n}\right|}\beta_{\overline{k}}({\cal K}_{n})\right]}_{{\rm Interference\leavevmode\nobreak\ of\leavevmode\nobreak\ NOMA\leavevmode\nobreak\ corresponding\leavevmode\nobreak\ to\leavevmode\nobreak\ the\leavevmode\nobreak\ other\leavevmode\nobreak\ users\leavevmode\nobreak\ in\leavevmode\nobreak\ }{\it{\cal I}_{k}}}+\underbrace{\left(\sum_{\tilde{n}\in{\cal I}_{k}^{\tilde{n}}}h_{{k},\tilde{n}}\left(\theta_{k,\tilde{n}},\psi_{k,\tilde{n}}\middle|\Psi\right)\right)^{2}}_{{\rm Interference\leavevmode\nobreak\ from\leavevmode\nobreak\ the\leavevmode\nobreak\ other\leavevmode\nobreak\ LEDs}}+\frac{\sigma_{N}^{2}}{P_{s}}}$ (17)
In this work, the transmission sum rate among the legitimate users and the secrecy sum rate are defined as [44]
$R_{T}$ $=\sum_{n=1}^{N}\sum_{k=1}^{K}R_{D_{k}}\left(\theta_{k,n},\psi_{k,n}\right),$ (18)
$R_{S}$ $=\sum_{n=1}^{N}\sum_{k=1}^{K}\left[R_{D_{k}}\left(\theta_{k,n},\psi_{k,n}\right)-R_{E,D_{k}}\left(\theta_{E,n},\psi_{E,n}\right)\right]^{+},$ (19)
respectively.
IV Transmission Strategies and Power Allocation Schemes
In this section, we introduce the three transmission strategies and two power allocation schemes investigated in this work.
IV-A Transmission Strategies
We first introduce the conventional broadcasting strategy used in [45]. Then, we propose two new strategies referred to as simple and smart LED linking strategies, depending on the various locating scenarios of users in order to improve their transmission performance.
IV-A1 Broadcasting Strategy
In this transmission strategy, all the LEDs transmit the same superposed signal by NOMA [45]. The transmitted signal by $S_{n}$, denoted by $x_{n}^{\rm b}$ for this strategy, is similar to Eq. 12 with ${\cal K}_{n}={\cal K}$, which can be expressed as
${x}^{\rm b}_{n}=\sum_{k\in\mathcal{K}}\sqrt{P_{s}\beta_{k}({\cal K})}s_{n,k}+I_{\mathrm{DC}}.$ (20)
IV-A2 Simple LED Linking Strategy
In this strategy, the system will transmit signals based on the estimated location of legitimate users. The LED will transmit signals to users located in its coverage area. In case there exist more than one user in the coverage of the LED, the signal will be transmitted by the multiple access scheme, i.e., NOMA. Hence, the transmitted signal by $S_{n}$, denoted by $x_{n}^{\rm smp}$, can be expressed as
${x}^{\mathrm{smp}}_{n}=\begin{cases}\sum_{k\in\mathcal{K}_{n}}\sqrt{\beta_{k}({\cal K}_{n})P_{s}}s_{n,k}+I_{\mathrm{DC}},&{\cal K}_{n}\neq\varnothing,\\ I_{\mathrm{DC}},&\text{otherwise}.\end{cases}$ (21)
A disadvantage of this method is that the devices located in the overlapping area may be interfered with the signals sent from different LED sources, and the interference causes some rate loss in terms of transmission performance. In order to overcome such a rate loss, we further introduce a novel smart LED linking strategy.
IV-A3 Smart LED Linking Strategy
Algorithm 1 Smart LED linking strategy
1:$D_{k}(x_{D_{k}},y_{D_{k}},z_{D_{k}})$ device coordinates, $S_{n}(x_{S_{n}},y_{S_{n}},Z)$ LED coordinates, $\theta_{1/2}$ half illuminance angle.
2:Group set of LEDs ${{\mathbb{L}}}$
3:Initialize: $r\leftarrow(Z-z_{D_{k}})\tan\theta_{1/2}$, ${{\mathbb{L}}}\leftarrow\varnothing$
4:for $n=1:N$ do
5:     ${\cal K}_{n}\leftarrow\varnothing$, ${\cal M}\leftarrow\varnothing$, ${\cal L}_{n}\leftarrow\varnothing$
6:     ${{\mathbb{L}}}\leftarrow{{\mathbb{L}}}\cup\{{\cal L}_{n}\}$triangleright Generate a new set for the $n$th LED
7:     for $k=1:N$ do
8:         if $(x_{D_{k}}-x_{S_{n}})^{2}+(y_{D_{k}}-y_{S_{n}})^{2}\leq r^{2}$ then $\triangleright$ $D_{k}$ in the coverage area of $S_{n}$
9:              ${\cal K}_{n}\leftarrow{\cal K}_{n}\cup\{k\}$
10:         end if
11:         $k\leftarrow k+1$
12:     end for
13:     for $l=1:n$ do
14:         if ${\cal L}_{l}\cap{\cal K}_{n}\neq\varnothing$ then $\triangleright$ Check any user indices duplicated between ${\cal S}_{l}$ and ${\cal S}_{n}$
15:              ${\cal M}\leftarrow{\cal M}\cup\{l\}$, ${\cal K}_{n}\leftarrow{\cal L}_{l}\cup{\cal K}_{n}$triangleright Record LED index to ${\cal M}$, and merge ${\cal L}_{l}$ to ${\cal K}_{n}$
16:         end if
17:     end for
18:     for $m\in{\cal M}$ do $\triangleright$ Overwrite ${\cal L}_{m}$ to ${\cal K}_{n}$ after the previous loop
19:         ${\cal L}_{m}\leftarrow{\cal K}_{n}$
20:     end for
21:     ${\cal L}_{n}\leftarrow{\cal K}_{n}$triangleright Record ${\cal K}_{n}$ to ${\cal L}_{n}$
22:end for
23:return ${{\mathbb{L}}}$
Since it may be difficult for PDs to detect the interfered symbol in the simple LED linking strategy, we propose the smart LED linking strategy in order to avoid receiving the interfering symbols. In this strategy, the system will first determine if there is any user in the overlapping coverage area. If so, the system will synchronize the two LEDs covering the overlapping area. These linked LEDs will transmit the same signal independently, according to the needs of users in that region. This strategy is described by Algorithm 1, where we denote the set of group sets of LEDs as ${\mathbb{L}}$, and the transmitted signal in the $n$th group set ${\cal L}_{n}$ (${\cal L}_{n}\in{\mathbb{L}}$) by $S_{n}$. Similar to the simple LED linking, the transmitted signal, denoted by $x_{n}^{\rm smt}$, is expressed as
$x^{\mathrm{smt}}_{n}=\begin{cases}\sum_{l\in{\cal L}_{n}}\sqrt{\beta_{l}({\cal L}_{n})P_{s}}s_{n,l}+I_{\mathrm{DC}},&{\cal L}_{n}\neq\varnothing,\\ I_{\mathrm{DC}},&\text{otherwise}.\end{cases}$ (22)
IV-B Power Allocation Schemes Based on Maximum Sum Rate
We introduce two distinct power allocation schemes within the context of NOMA in this work. The first scheme is the conventional fixed power allocation scheme [46]. The second, as our novel contribution, is the maximum sum rate-based scheme, which optimizes power distribution by estimating the maximum sum rate, thereby significantly enhancing the performance and efficiency of NOMA systems.
IV-B1 Fixed Power Allocation Scheme
In this scheme, we assume that the power ratio among different users is fixed. In SIC, the receiver will detect the signal with stronger power first. Therefore, one of the power ratio parameters of NOMA $\beta_{k}^{\mathrm{fix}}({\cal K}_{n})$ for $k\in{\cal K}_{n}$ can be defined as [46]
$\beta_{k}^{\mathrm{fix}}({\cal K}_{n})\triangleq\begin{cases}\zeta(1-\zeta)^{k-1},&\leavevmode\nobreak\ \mbox{if}\leavevmode\nobreak\ \leavevmode\nobreak\ k<|{\cal K}_{n}|,\\ (1-\zeta)^{k-1},&\leavevmode\nobreak\ \mbox{if}\leavevmode\nobreak\ \leavevmode\nobreak\ k=|{\cal K}_{n}|,\end{cases}$ (23)
where $\zeta\in(0.5,1]$ is the fixed power ratio.
IV-B2 Maximum Sum Rate-based Scheme
In this scheme, the LED is able to optimize the power allocation in order to maximize the sum rate only by the positions of users. By defining the power ratio parameter of NOMA in this scheme as $\beta_{k}^{\mathrm{opt}}({\cal K}_{n})$, the corresponding optimization problem of maximizing the sum of the optimized transmission rate $\widehat{R}_{k}$ is formulated as follows [47]:
maximize $\sum_{k\in\mathcal{K}}\widehat{R}_{k}\left(\overline{\lambda}_{k},\tilde{\omega}_{k,n},\beta_{k}^{\mathrm{opt}}({\cal K}_{n})\right)$ (24)
$\mathrm{subject\leavevmode\nobreak\ to}\quad$ $\sum_{k\in\mathcal{K}}\beta_{k}^{\mathrm{opt}}({\cal K}_{n})\leq 1,$ (24a)
$\beta_{\hat{k}}^{\mathrm{opt}}({\cal K}_{n})\geq\beta_{\hat{k}+1}^{\mathrm{opt}}({\cal K}_{n}),$ (24b)
$\beta_{k}^{\mathrm{opt}}({\cal K}_{n})\geq 0$ (24c)
for $\hat{k}\in\{1,2,\ldots,K-1\}$. In this optimization problem, the radiation angle $\overline{\lambda}_{k}$ is set as the expected value of $\lambda_{k}$ defined in Section II-B1 and the azimuth angle $\tilde{\omega}_{k,n}$ of $D_{k}$ is set as the argument of the maximum sum rate. As a result, the formula of the optimization problem can be rewritten as
$\sum_{k\in\mathcal{K}}\widehat{R}_{k}\left(\overline{\lambda}_{k},\tilde{\omega}_{k,n},\beta_{k}^{\mathrm{opt}}({\cal K}_{n})\right)$
$=\sum_{k\in\mathcal{K}}\frac{1}{2}\log_{2}\left(1+\widehat{\gamma}_{k}\left(\overline{\lambda}_{k},\tilde{\omega}_{k,n},\beta_{k}^{\mathrm{opt}}({\cal K}_{n}),{\cal K}_{n}\right)\right),$ (25)
where $\widehat{\gamma}_{k}\left(\overline{\lambda}_{k},\tilde{\omega}_{k,n},\beta_{k}^{\mathrm{opt}}({\cal K}_{n}),{\cal K}_{n}\right)$ is defined in Eq. 26 located at the top of this page
$\widehat{\gamma}_{k}\left(\overline{\lambda}_{k},\tilde{\omega}_{k,n},\beta_{k}^{\mathrm{opt}}({\cal K}_{n}),{\cal K}_{n}\right)$
=$=\left\{\displaystyle\overbrace{\max_{{\cal K}_{n}}\left[\left(\sum_{n\in{\cal I}_{k}}\hat{h}_{{k},n}\left(\theta_{k,n},\hat{\psi}_{k,n}(\overline{\lambda}_{k},\tilde{\omega}_{k,n})\middle|\Psi\right)\right)^{2}\beta_{k}({\cal K}_{n})\right]}^{{\rm Decoded\leavevmode\nobreak\ information\leavevmode\nobreak\ of\leavevmode\nobreak\ }{\it k}{\rm th\leavevmode\nobreak\ user\leavevmode\nobreak\ from\leavevmode\nobreak\ }{\it{\cal I}_{k}}}\right\}$
$\cdot\left\{\displaystyle\underbrace{\max_{{\cal K}_{n}}\left[\left(\sum_{n\in{\cal I}_{k}}\hat{h}_{{k},n}\left(\theta_{k,n},\hat{\psi}_{k,n}(\overline{\lambda}_{k},\tilde{\omega}_{k,n})\middle|\Psi\right)\right)^{2}\sum_{\overline{k}={k+1}}^{\left|{\cal K}_{n}\right|}\beta_{\overline{k}}({\cal K}_{n})\right]}_{{\rm Interference\leavevmode\nobreak\ of\leavevmode\nobreak\ NOMA\leavevmode\nobreak\ from\leavevmode\nobreak\ other\leavevmode\nobreak\ users\leavevmode\nobreak\ in\leavevmode\nobreak\ }{\it{\cal I}_{k}}}+\underbrace{\left(\sum_{\tilde{n}\in{\cal I}_{k}^{\tilde{n}}}\hat{h}_{{k},\tilde{n}}\left(\theta_{k,\tilde{n}},\hat{\psi}_{k,\tilde{n}}(\overline{\lambda}_{k},\tilde{\omega}_{k,\tilde{n}})\middle|\Psi\right)\right)^{2}}_{{\rm Interference\leavevmode\nobreak\ from\leavevmode\nobreak\ the\leavevmode\nobreak\ other\leavevmode\nobreak\ LEDs}}+\frac{\sigma_{N}^{2}}{P_{s}}\right\}^{-1}$ (26)
with the estimated channel gain $\hat{h}_{{k},n}\left(\theta_{k,n},\hat{\psi}_{k,n}(\overline{\lambda}_{k},\tilde{\omega}_{k,n})\middle|\Psi\right)$ formulated as
$\hat{h}_{{k},n}$ $\left(\theta_{k,n},\hat{\psi}_{k,n}(\overline{\lambda}_{k},\tilde{\omega}_{k,n})\middle|\Psi\right)$
$=A\frac{(m+1)R_{\mathrm{PD}}}{2\pi}\cos^{m}\left({\theta}_{k,n}\right)\frac{\cos\left(\hat{\psi}_{k,n}(\overline{\lambda}_{k},\tilde{\omega}_{k,n})\right)}{d_{D_{k},n}^{2}}$
$\quad\quad\cdot g\left(\hat{\psi}_{k,n}(\overline{\lambda}_{k},\tilde{\omega}_{k,n})\middle|\Psi\right),$ (27)
$\overline{\lambda}_{k}$ $={\mathbb{E}}\left[{\lambda_{k}}\right],$ (28)
$\tilde{\omega}_{k,n}$ $=\mathop{\arg\max}_{\omega_{k,n}\in[0,2\pi)}\left(\hat{h}_{{k},n}\left(\theta_{k,n},\hat{\psi}_{k,n}(\overline{\lambda}_{k},\omega_{k,n})\middle|\Psi\right)\right).$ (29)
Note that $\hat{\psi}_{k,n}(\overline{\lambda}_{k},\tilde{\omega}_{k,n})$ can be derived by substituting Eqs. 28 and 29 into Section II-B1. Even though Eq. 24 is a non-linear constrained optimization problem, one can solve it numerically by JuMP package with Ipopt optimizer [48, 49]666 The case of the indicator function introduces non-convexity into the optimization problem defined in Eq. 24, specifically with regard to the channel gain $\hat{h}_{k,n}$. This non-convex nature poses a challenge for applying convex optimization or approximation methods to address the problem. However, it is important to note that the focus of the current study does not extend to exploring convex optimization for this particular challenge. Instead, our interest lies in establishing the maximum on the transmission and secrecy performance. Despite the drawback of high complexity, the chosen optimizer is sufficient for achieving our objectives. .
V Simulation Results and Discussion
In this section, we first describe the setups of the numerical environment in our proposed VLC system model with the mathematical body blockage. Additionally, we evaluate the transmission sum rate of the users as well as the secrecy sum rate. We also investigate the performance difference among transmission strategies and the power allocation schemes of NOMA.
V-A Parameter Settings for Simulations
TABLE III: Simulation parameters
Room configuration
Length $L$$\times$ Width $W$ 40 $\times$ 40 $\mathrm{m}^{2}$
Height of the room $Z$ 3.98 $\mathrm{m}$
Number of LED arrays 23
Number of users $K$ 6
The side length of the equilateral triangle of LED sources $l$ $9.6\leavevmode\nobreak$m
LED electrical and optical characteristics
Average optical power per LED 0.25 W
Half-intensity angle $\theta_{\frac{1}{2}}$ $70^{\circ}$
Power ratio of NOMA $\zeta$ 0.6
Photodiode characteristics
Received fiend $\Psi$ $60^{\circ}$
Physical area $A$ $\text{cm}^{2}$
Noise power $\sigma_{N}^{2}$98.35 dBm
Reflective index $\eta$ 1.5
Height ratio parameter of the device $\nu$ 0.75
Parameters related to the bodies of users and eavesdropper
Mean of the height of the users and eavesdropper $H$ 1.6 $\mathrm{m}$
Standard deviation of the height of the users and the eavesdropper $\sigma_{H}$ 0.1333 $\mathrm{m}$
Distance from the center of the top base of cylinder to the device $l_{d}$ 0.4 m
Radius of bodies $r$ 0.2 m
Algorithm 2 Random height generation algorithm.
1:The average body height $H$ and its standard deviation $\sigma_{H}$.
2:${\sf h}_{D_{i}}$ with $i\in{\cal K}^{\prime}$.
3:for all $i\in{\cal K}^{\prime}$ do
4:     while True do
5:         Generate ${\sf h}_{D_{i}}$ with ${\sf h}_{D_{i}}\sim{\cal N}(H,\sigma_{H}^{2})$.
6:         if ${\sf h}_{i}$ > 0 then
7:              return ${\sf h}_{D_{i}}$.
8:         end if
9:     end while
10:end for
Figure 6: The arrangement of LED sources (blue dots represent the positions of LEDs and unit of horizontal and vertical axes are measured by meters).
TABLE IV: The Cartesian XOY coordinates of users.
Scenario 1 Scenario 2 Scenario 3
$D_{1}$ $(6,6)$ $(13,16)$ $(10.6,14.5)$
$D_{2}$ $(34,6)$ $(20,12)$ $(15.3,22.7)$
$D_{3}$ $(34,34)$ $(27,16)$ $(5.9,22.7)$
$D_{4}$ $(6,34)$ $(27,24)$ $(34.1,20)$
$D_{5}$ $(20,10)$ $(20,28)$ $(34.1,32.2)$
$D_{6}$ $(20,30)$ $(13,24)$ $(34.1,7.8)$
(a) Scenario 1
(b) Scenario 2
(c) Scenario 3
Figure 7: Illustration of the location of PDs.
The parameters for the numerical simulations are listed in Table III, including LED and PD settings, as well as the body blockages [33, 28]. Furthermore, in order to reduce the repetitive coverage by different LED units, we consider a new arrangement of LEDs where each LED is placed in the vertexes of equilateral triangles. This arrangement is illustrated in Fig. 6. The side length of the triangle satisfies the following inequality:
$\frac{l}{2}\frac{1}{\cos\frac{\pi}{6}}\leq(z-z_{D_{k}})\tan\theta_{1/2}.$ (30)
Therefore, we have
$l\leq\sqrt{3}(z-z_{D_{k}})\tan\theta_{1/2}.$ (31)
We set the location of the first LED at $(20,20)$ in the ceiling, and the remaining LEDs are assigned to the blue points (cf. Fig. 6), where the coordinates of any three adjacent LEDs form an equilateral triangle with side length $l$. The geometric coordinates of users are set as in Table IV with variance $z_{D_{k}}$ in each simulation loop (also illustrated in Fig. 7); Scenario 1 corresponds to the case where all the users are sparsely located in the room (cf. Fig. 7(a)), Scenario 2 corresponds to the case where all the users are closely located in the room, and finally (cf. Fig. 7(b)), Scenario 3 corresponds to the case of a hybrid combination of the two scenarios (cf. Fig. 7(c)). The Monte Carlo simulation method is employed, with $10^{4}$ iterations determined by the trade-off between simulation accuracy and computational complexity. Also, in this work, the height of each user ${\sf h}_{D_{k}}$ as well as the eavesdropper ${\sf h}_{D_{E}}$ is generated by the Algorithm 2. Therefore, in each simulation loop, we first generate the heights of each user and the eavesdropper. Afterward, we calculate the transmission rate for that loop.
V-B Transmission Performance
Figure 8: Transmission power versus transmission sum rate with fixed power allocation scheme while the coordinates of the eavesdropper are uniformly distributed (solid lines with circle markers are the graphs of the broadcasting strategy, dash lines with triangle markers are the graphs of the simple LED linking strategy, and dot lines with square markers are the graphs of the smart LED linking strategy).
In Fig. 8, we investigate the transmission sum rate of the three aforementioned transmission scenarios with the proposed strategies and fixed power allocation scheme when $x_{E}$ and $y_{E}$ are uniformly distributed from 1 m to 39 m (i.e., in the distributed area), where the cylinder bodies may block the lights received by the users. From this figure, we can observe that our proposed simple LED linking and smart LED linking strategies outperform the conventional broadcasting strategy. For a comparison of the performance of the two LED linking strategies, the results depend on how users are distributed. In Scenarios 1 and 2, the simple LED linking strategy outperforms the smart LED linking strategy, whereas the smart LED linking strategy outperforms the simple LED linking strategy in Scenario 3. From Fig. 8, we also observe that this performance difference is not affected by the transmission power due to the fact that the location of the users is fixed. Apparently, the distribution of the users affects the transmission performance of the system.
Figure 9: Transmission power versus transmission sum rate with the fixed power allocation scheme and maximum sum rate-based scheme.
We next analyze the transmission performance of our model under the three strategies, considering both the maximum sum rate-based scheme and the fixed power allocation scheme. The simulation results are depicted in Fig. 9. From the figure, we can observe that in Fig. 9(a), the transmission performance in the broadcasting strategy for Scenario 2 is poor, and this is due to the fact that the maximum sum rate-based scheme is only based on the expected value of the radiation and azimuth angles, and the power allocation calculation process will cause severe errors when the users are densely located. Besides, the maximum sum rate-based scheme works well with the two LED linking strategies. As shown in Figs. 9(b) and 9(c), the values of the transmission sum rates for this scheme are basically higher than that for the fixed power allocation scheme based on the sum rate maximization of the legitimate users. Furthermore, we can observe that in Fig. 9(c), the maximum sum rate-based scheme in Scenario 1 and Scenario 3 does not significantly improve the performance due to the fact that the locations of users are sparser than those in Scenario 2, i.e., each ${\cal L}_{n}$ contains only a single or two users. From these observations, we may conclude that under a realistic VLC system where the light path could be blocked, the newly proposed strategies (simple and smart LED linking strategies) provide better gain for the transmission sum rate compared to the broadcasting strategy.
Figure 10: Transmission power versus transmission sum rate with fixed power allocation scheme in the proposed LED arrangement and the LED arrangement in [28] with the same total transmission power.
Additionally, in Fig. 10, we compare the transmission sum rate between our proposed LED arrangement and the squared LED arrangement [28] with the same sidelength and total transmission power. From this figure, we observe that in Scenario 3, our proposed scenario outperforms the LED arrangement in [28], but in Scenarios 1 and 2, the transmission sum rate for our proposed arrangement is inferior to that of the arrangement method in [28]. This can be attributed to the fact that the squared LED arrangement increases the overlapping area of two geometrically adjacent LEDs, resulting in an increased illuminance for each user. If the two adjacent different sources transmit the same symbol simultaneously, the transmission performance can be improved. It is worth noting that the transmission performance of the LED arrangements is also significantly influenced by the locations of users, which requires further investigation.
Figure 11: NOMA power ratio $\zeta$ versus transmission sum rate with different LED arrangements and transmission power $P_{s}=0.6$ W.
Conversely, to observe the effects of the NOMA power ratio $\zeta$, the transmission performance is plotted with respect to $\zeta$ in Fig. 11. Notably, the transmission sum rates for the broadcasting strategy and smart LED linking strategy remain constant, while the simple LED linking strategy experiences a significant decrease as $\zeta$ increases. The reason for this phenomenon is that, in the simple LED linking strategy, users may receive signals from two different adjacent sources based on their location. The power of the interference decreases with the increase of $\zeta$.
V-C Secrecy Performance
(a) Scenario 1
(b) Scenario 2
(c) Scenario 3
Figure 12: Three-dimensional plots of XOY location of the eavesdropper versus secrecy sum rate with the fixed power allocation scheme in the broadcasting strategy.
(a) Scenario 1
(b) Scenario 2
(c) Scenario 3
Figure 13: Three-dimensional plots of XOY location of the eavesdropper versus secrecy sum rate with the fixed power allocation scheme in the simple LED linking strategy.
(a) Scenario 1
(b) Scenario 2
(c) Scenario 3
Figure 14: Three-dimensional plots of XOY location of the eavesdropper versus secrecy sum rate with the fixed power allocation scheme in the smart LED linking strategy.
The secrecy sum rate and the location of the eavesdropper of our system model are examined with the three transmission strategies under the fixed power allocation scheme. The simulation results for the broadcasting, simple LED linking, and smart LED linking strategies are shown in Figs. 12, 13, and 14, respectively. From these figures, it is evident that the broadcasting strategy has the lowest secrecy sum rate. The secrecy performances achieved by the simple as well as smart LED linking strategies rely on the distribution of the users. More precisely, for Scenario 2, corresponding to the case where the users are in a small area with high concentration, the simple LED linking strategy provides better secrecy performance than the smart one (c.f., Figs. 13(b) and 14(b)). The reason for this behavior is that the LEDs have the grouping process before transmitting data to users in the smart LED linking strategy in order to bypass signal interference in the overlapping regions. When all the users are densely located, they may be assigned to the same group. Therefore, only a subset of the entire LEDs is selected to transmit signals to the legitimate users, and the remaining LEDs make no contribution to increasing the secrecy sum rate. In different circumstances, if some users are sparsely distributed as in Scenario 3, the smart LED linking strategy provides a higher secrecy sum rate than the simple one, which can easily be seen by comparing Fig. 13(c) and Fig. 14(c). This is because the eavesdropper cannot wiretap the messages of all the users, and the signal interference from different LEDs is also mitigated, resulting in boosting the secrecy sum rate of VLC systems.
VI Conclusion
In this work, we investigated the transmission and the secrecy performances of the indoor VLC networks for mobile devices equipped with a single PD under NOMA transmission. Specifically, we proposed a geometric body blockage model with a novel LED arrangement and evaluated our proposed model with three different transmission strategies, i.e., broadcasting, simple LED linking, and smart LED linking strategies. As a result, compared to the conventional broadcasting strategy, our proposed simple and smart LED linking strategies provided significant improvements in terms of transmission and secrecy performances in the simulation results.
For future work, the LED arrangement for the different users’ locations needs further investigation, and it may be of significant interest to derive explicit theoretical bounds of the transmission sum rate and the secrecy sum rate for the body blockage model.
Appendix A Proof of Lemma 1
With reference to Fig. 5, a vector on the plane $\mathcal{W}$ and its normal vector ${\bm{v}}_{s}$ are orthogonal and thus
$\overrightarrow{NP}\cdot{\bm{v}}_{s}=0.$ (A.1)
The straight line that has the directional vector ${\bm{v}}_{t}$ and pass through the point $M$ can be expressed as
$\overrightarrow{OP}=\overrightarrow{OM}+\alpha{\bm{v}}_{t},$ (A.2)
where $O$ denotes the origin and $\alpha\in{\mathbb{R}}$.
Substituting Eq. A.2 into Eq. A.1, we have
$((\overrightarrow{OM}+\alpha{\bm{v}}_{t})-\overrightarrow{ON})\cdot{\bm{v}}_{s}=0,$
and thus
$\alpha=\frac{(\overrightarrow{ON}-\overrightarrow{OM})\cdot{\bm{v}}_{s}}{{\bm{v}}_{t}\cdot{\bm{v}}_{s}}=\frac{\overrightarrow{MN}\cdot{\bm{v}}_{s}}{{\bm{v}}_{s}\cdot{\bm{v}}_{t}}.$ (A.3)
Now substitution of Eq. A.3 into Eq. A.2 yields that
$\overrightarrow{OP}=\overrightarrow{OM}+\frac{\overrightarrow{MN}\cdot{\bm{v}}_{s}}{{\bm{v}}_{s}\cdot{\bm{v}}_{{t}}}{\bm{v}}_{{t}},$
i.e.,
$\overrightarrow{MP}=\frac{\overrightarrow{MN}\cdot{\bm{v}}_{s}}{{\bm{v}}_{s}\cdot{\bm{v}}_{{t}}}{\bm{v}}_{{t}},$ (A.4)
which completes the proof of Lemma 1. ∎